OpenStudy (anonymous):

Rewrite cos 3x with only sin x and cos x

OpenStudy (sidsiddhartha):

start with this Cos(A+B) = CosA.CosB - SinA.SinB here cos(x+2x) = cosx.cos2x - sinx.sin2x

OpenStudy (sidsiddhartha):

now use cos(2x)=(2cos^2x - 1) and sin2x=2sinx.cosx